For the hockey tournament, two buses leave 20 minutes apart and travel along the same road. The first bus travels at 66km/h and the later bus, 77km/h. How long will it take before the later bus catches up?
(I know the answer : 2 hours, but I don't know how to show my work. I know it involves something like how much the first bus travelled before the later bus. { 11km })

Answer by checkley79(3341)

D=RT
D=66T FOR THE FIRST BUS.
D=77(T-20/60) OR
D=77(T-.33) FOR THE SECOND BUS.
SINCE THE 2 DISTANCES ARE EQUAL THUS:
66T=77(T-.33)
66T=77T-25.41
66T-77T=-25.41
-11T=-25.41
T=-25.41/-11
T=2.31 HOURS.
PROOF:
66*2.31=77(2.31-.33)
152.46=77*1.98
152.46=152.46
#########################################
IF THE ANSWER IS 2 HOURS THEN:
66*2=77(2-20/60)
132=77(2-.33)
132=77*1.67
132=128.33 DOES NOT COMPUTE.